# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 6 Conclusion
We start with partitioning the unobserved responses vector into $k$ blocks, such that $\textbf{y}_{u}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$. Using proposals from $p(\textbf{y}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u}^{(-j)})$, the Algorithm S2 outlines the MCMC steps for sampling the missing values one block at a time, for $j=1,..,k$, where $\textbf{y}_{u_{j}}$ is the updated block and $\textbf{y}_{u}^{(-j)}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{j-1}}^{\top},\textbf{y}_{u_{j+1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$ is the remaining blocks. The complete response vector y can now be written as $\textbf{y}=({\textbf{y}^{\top}_{s_{j}}},\textbf{y}_{u_{j}}^{\top})^{\top}$, where ${\textbf{y}_{s_{j}}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top})^{\top}$. Based on this partitioning of y, the following partitioned vector and matrices are defined:
@@
where $\textbf{r}_{s_{j}}$ is the corresponding sub-vector of r for the observed responses ($\textbf{y}_{o}$), and the unobserved responses that are not in the $j^{th}$ block (${\textbf{y}_{u}^{(-j)}}$). i.e. $\textbf{r}_{s_{j}}=({{\textbf{r}}}^{\top}_{o},{{\textbf{r}}^{\top}_{u^{(-j)}}})^{\top}$ and ${\textbf{r}}_{u_{j}}$ is the corresponding sub vector of r for $j^{th}$ block of unobserved responses, $\textbf{y}_{u_{j}}$. Similarly, $\textbf{X}_{s_{j}}$ and $\textbf{X}_{u_{j}}$ are represent the sub-matrices of X, and $\textbf{M}_{{s_{j}}{s_{j}}}$, $\textbf{M}_{{s_{j}}{u_{j}}}$, $\textbf{M}_{{u_{j}}{s_{j}}}$, and $\textbf{M}_{{u_{j}}{u_{j}}}$ are sub-matrices of M.
@@
6:        Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{i-1},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top},{\widetilde{\textbf{y}}_{u_{j}}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top},{\textbf{y}_{u_{j},{i-1}}}^{\top})^{\top}$
@@
In step 4 of the Algorithm S2, proposals from the conditional distributions, $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{s_{j}})$ needs to be generated. For the SEM-Gau, and SEM-t, this distribution follows a multivariate Gaussian distribution with the mean $\textbf{X}_{u_{j}}\boldsymbol{\beta}-\textbf{M}_{{u_{j}}{u_{j}}}^{-1}\textbf{M}_{{u_{j}}{s_{j}}}\textbf{r}_{s_{j}}$ and the covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}_{{u_{j}}{u_{j}}}^{-1}$.
@@
S4.3 Expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$
This section presents the derivatives $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, which are required to compute the gradients of the reparameterised ELBO, defined in Equation (17) of the main paper for different SEMs with missing data.
S4.3.1 Derivatives for SEM-Gau
For SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S48)
where $\textbf{z}_{i}=({\textbf{x}^{*}_{i}},y_{i})$ is the vector containing the $i^{th}$ row vector of matrix $\textbf{X}^{*}$, and the $i^{th}$ element of the vector y, see Section 2.3 of the main paper.
S4.3.2 Derivatives for SEM-t
For SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\nu,\rho)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\nu$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S50)
S4.3.3 Derivatives for YJ-SEM-Gau
For YJ-SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S51)
S4.3.4 Derivatives for YJ-SEM-t
For YJ-SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$. The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~{}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~{}p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~{}p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\phi})+\textrm{log}~{}p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section 4 of the main paper, and put prior on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{(\textbf{x}^{*}_{i}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~{}p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}},\end{split}$ (S52)
S5 Bayesian model comparison of spatial error models
@@
This section presents expressions for the term log $p(\mathbf{y}\mid\boldsymbol{\phi})$, which are used to compute the Deviance Information Criterion (DIC) (Celeux et al.,, 2006) for the various SEMs discussed in Section 3.3 of the main paper.
Note that for SEM-Gau, and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, meaning for these two models $p(\mathbf{y}\mid\boldsymbol{\phi})=p(\mathbf{y}\mid\boldsymbol{\xi})$. For the SEM-Gau model, $p(\mathbf{y}\mid\boldsymbol{\xi})$ is Gaussian; see Section 2.1 of the main paper. The log-likelihood is given by:
@@
where M, and r are given in Table S1.
For the YJ-SEM-Gau, $p(\mathbf{y}\mid\boldsymbol{\phi})=p(\mathbf{y}\mid\boldsymbol{\xi})$ is derived in Section 2.2 of the main paper, and the log-likelihood is given by:
@@
For the SEM-t model,
[…]
where $p(\mathbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})$ is multivariate Gaussian, with its mean vector and covariance matrix specified in Table S1. Each $p({\tau}_{i}\mid\boldsymbol{\phi})$ (for $i=1,\ldots,n$) follows an inverse gamma distribution with both shape and scale parameters equal to $\nu/2$; see Section 2.1 of the main paper. This hierarchical representation implies that the marginal distribution $p(\mathbf{y}\mid\boldsymbol{\phi})$ is a multivariate Student’s $t$-distribution with mean vector $\textbf{X}\boldsymbol{\beta}$, scale matrix $\sigma^{2}_{\boldsymbol{e}}(\mathbf{A}^{\top}\mathbf{A})^{-1}$, and degrees of freedom $\nu$. The log-likelihood is given by:
@@
where $\textbf{r}=\textbf{y}-\textbf{X}\boldsymbol{\beta}$.
Similarly, logarithm of $p(\textbf{y}\mid\boldsymbol{\phi})$ for YJ-SEM-t can be derived as
@@
S6 Simulation study with full data SEMs
@@
This section fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models to simulated datasets 1 and 2 (without missing data) in Section 4.2 of the main paper, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets.
@@
The initial values for the VB algorithms are set as follows: The variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE). The initial values for $\nu$ and $\gamma$ are set to 4 and 1, respectively. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors.
After some experimentation, we found that the models with Student’s $t$ errors required a relatively large number of iterations to converge. This increased computational demand is attributed to the added complexity of estimating both the latent variables $\boldsymbol{\tau}$ and the degrees of freedom parameter $\nu$. For simulated dataset 1, the VB algorithms were run for 10,000 iterations for both SEM-Gau and YJ-SEM-Gau, and 20,000 iterations for SEM-t and YJ-SEM-t. For simulated dataset 2, convergence was achieved after 10,000 iterations for SEM-Gau and YJ-SEM-Gau, and 40,000 iterations for SEM-t and YJ-SEM-t. Convergence is assessed visually by inspecting the variational mean plots over iterations, which are presented in Section S9.1. Then, we generate 10,000 draws of model parameters from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$.
@@
Table S4 shows the posterior means and the 95% credible intervals for some of the model parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models obtained using the Gaussian variational approximation method described in Section S2 applied to the simulated dataset 1. The table also includes the computation times (in seconds) for one VB iteration, as well as the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model, calculated using Equations (18) and (19) of the main paper, respectively. The plots of posterior densities of the parameters are provided in Section S7.2.1.
We begin by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and $\beta_{1}$), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are present in all four models. Accurate estimation of these parameters is crucial for interpreting the underlying spatial relationships and the effect of covariates on the response variable in SEMs. The posterior means for $\beta_{0}$, $\beta_{1}$ and $\rho$ are nearly identical for YJ-SEM-Gau and YJ-SEM-t, closely matching the true values of $-1$ for $\beta_{0}$, $1$ for $\beta_{1}$, and $0.8$ for $\rho$. In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values. The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau deviate considerably from the true value, while the posterior mean obtained from YJ-SEM-t is much closer to the true value, as expected. The models incorporating the YJ transformation, YJ-SEM-Gau and YJ-SEM-t, successfully recover the parameter $\gamma$. The degrees of freedom ($\nu$) estimated by the SEM-t and YJ-SEM-t models show slight deviations from the true value.
For the simulated dataset 1, the model fit assessment indicates that YJ-SEM-t is the best-performing model, as it achieves the lowest $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values among all four models. This outcome is expected, given that the dataset was simulated from YJ-SEM-t. The second-best model is YJ-SEM-Gau, which has the next lowest DIC values. Notably, the difference in DIC between YJ-SEM-t and YJ-SEM-Gau is minimal, suggesting that YJ-SEM-Gau provides a model fit nearly as good as YJ-SEM-t for this dataset.

### v2

@@ 9 Conclusion
We start with partitioning the unobserved responses vector into $k$ blocks, such that $\textbf{y}_{u}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$. Using proposals from $p(\textbf{y}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u}^{(-j)})$, Algorithm S2 outlines the MCMC steps for sampling the missing values one block at a time, for $j=1,..,k$, where $\textbf{y}_{u_{j}}$ is the updated block and $\textbf{y}_{u}^{(-j)}=(\textbf{y}_{u_{1}}^{\top},\dots,\textbf{y}_{u_{j-1}}^{\top},\textbf{y}_{u_{j+1}}^{\top},\dots,\textbf{y}_{u_{k}}^{\top})^{\top}$ is the remaining blocks. The complete response vector y can now be written as $\textbf{y}=({\textbf{y}^{\top}_{s_{j}}},\textbf{y}_{u_{j}}^{\top})^{\top}$, where ${\textbf{y}_{s_{j}}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u}^{(-j)}}^{\top})^{\top}$. Based on this partitioning of y, the following partitioned vector and matrices are defined:
@@
where $\textbf{r}_{s_{j}}$ is the corresponding sub-vector of r for the observed responses ($\textbf{y}_{o}$), and the unobserved responses that are not in the $j^{th}$ block ($\textbf{y}_{u}^{(-j)}$}). i.e. $\textbf{r}_{s_{j}}=({{\textbf{r}}}^{\top}_{o},{{\textbf{r}}^{\top}_{u^{(-j)}}})^{\top}$ and ${\textbf{r}}_{u_{j}}$ is the corresponding sub vector of r for $j^{th}$ block of unobserved responses, $\textbf{y}_{u_{j}}$. Similarly, $\textbf{X}_{s_{j}}$ and $\textbf{X}_{u_{j}}$ are represent the sub-matrices of X, and $\textbf{M}_{{s_{j}}{s_{j}}}$, $\textbf{M}_{{s_{j}}{u_{j}}}$, $\textbf{M}_{{u_{j}}{s_{j}}}$, and $\textbf{M}_{{u_{j}}{u_{j}}}$ are sub-matrices of M. Further, r and M for different SEMs are provided in Table S1.
@@
6:   Calculate $a=\text{min}\left(1,\frac{p(\textbf{m}\mid\widetilde{\textbf{y}},\boldsymbol{\psi}^{(t)})}{p(\textbf{m}\mid\textbf{y}_{i-1},\boldsymbol{\psi}^{(t)})}\right)$, where $\widetilde{\textbf{y}}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u,i-1}^{(-j)}}^{\top},{\widetilde{\textbf{y}}_{u_{j}}}^{\top})^{\top}$ and $\textbf{y}_{i-1}=(\textbf{y}_{o}^{\top},{\textbf{y}_{u,i-1}^{(-j)}}^{\top},{\textbf{y}_{u_{j},{i-1}}}^{\top})^{\top}$}
@@
In step 4 of Algorithm S2, proposals from the conditional distributions, $p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{o},\textbf{y}_{u,{i-1}}^{(-j)})={p(\widetilde{\textbf{y}}_{u_{j}}\mid\boldsymbol{\xi}^{(t)},\textbf{y}_{s_{j}})}$} needs to be generated. For the SEM-Gau, and SEM-t, this distribution follows a multivariate Gaussian distribution with the mean $\textbf{X}_{u_{j}}\boldsymbol{\beta}-\textbf{M}_{{u_{j}}{u_{j}}}^{-1}\textbf{M}_{{u_{j}}{s_{j}}}\textbf{r}_{s_{j}}$ and the covariance matrix $\sigma^{2}_{\textbf{e}}\textbf{M}_{{u_{j}}{u_{j}}}^{-1}$. Expressions of r and M for SEM-Gau and SEM-t are provided in Table S1, and details on the partitioning of r, X, and M are given in Equation (S54).
@@
S6 Bayesian model comparison for SEMs
@@
This section presents expressions for the term log $p(\textbf{y}\mid\boldsymbol{\phi})$ for different SEMs. These terms are used to compute the deviance information criterion (DIC) (Celeux et al.,, 2006) values discussed in Section 5 of the main paper. Note that for the SEM-Gau, and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, meaning for these two models $p(\textbf{y}\mid\boldsymbol{\phi})=p(\textbf{y}\mid\boldsymbol{\xi})$. For the SEM-Gau, $p(\textbf{y}\mid\boldsymbol{\xi})$ is multivariate Gaussian; see Section 2.1 of the main paper. Corresponding log-likelihood is given by:
@@
where M and r are given in Table S1.
For the YJ-SEM-Gau, $p(\textbf{y}\mid\boldsymbol{\phi})=p(\textbf{y}\mid\boldsymbol{\xi})$ is derived in Section 2.2 of the main paper. Corresponding log-likelihood is given by:
@@
For the SEM-t,
$p(\textbf{y}\mid\boldsymbol{\phi})=\int_{0}^{\infty}\ldots\int_{0}^{\infty}p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p({\tau}_{1}\mid\boldsymbol{\phi})\ldots p({\tau}_{n}\mid\boldsymbol{\phi})d{\tau}_{1}\ldots d{\tau}_{n}$ (S57)
where $p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})=p(\textbf{y}\mid\boldsymbol{\xi})$ is multivariate Gaussian, with its mean vector and covariance matrix specified in Table S1. Each $p({\tau}_{i}\mid\boldsymbol{\phi})$ (for $i=1,\ldots,n$) follows an inverse gamma distribution where both the shape parameter and the rate parameter are set to $\nu/2$; see Section 2.1 of the main paper. This hierarchical representation implies that the marginal distribution $p(\mathbf{y}\mid\boldsymbol{\phi})$ is a multivariate Student’s $t$-distribution with mean vector $\textbf{X}\boldsymbol{\beta}$, scale matrix $\sigma^{2}_{\boldsymbol{e}}(\mathbf{A}^{\top}\mathbf{A})^{-1}$, and degrees of freedom $\nu$. Then, the corresponding log-likelihood is given by:
@@
where $\textbf{r}=\textbf{y}-\textbf{X}\boldsymbol{\beta}$, and $\Gamma(\cdot)$ is the gamma function.
Similarly, the logarithm of $p(\textbf{y}\mid\boldsymbol{\phi})$ for YJ-SEM-t can be derived as
@@
S7 Additional figures for simulation study: Assessing the accuracy of VB methods in Section 6
Section 6 of the main paper compares the posterior density estimates of the model parameters for YJ-SEM-Gau obtained using VB (and HVB when missing values are present) and HMC (Neal,, 2011), based on a simulated dataset analysed both with and without missing values. To generate the data, the model parameters are set as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between 3-3 and $3$ (excluding 0). The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autoregressive parameter is $\rho=0.8$. The Yeo-Johnson transformation parameter is set to $\gamma=1.25$, which introduces moderate left skewness in the simulated response variable. The kernel density plot of the simulated response variable is provided in Figure S1. As expected, the density is moderately skewed to the left.
Figure S1: Kernel density plot of the simulated response variable generated using the YJ-SEM-Gau with $\gamma=1.25$ in the simulation study described in Section 6 of the main paper.
@@
S8 Simulation study: Assessing the accuracy and robustness of proposed SEMs (full data case)
@@
This section fits SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t to simulated datasets 1 and 2 (without missing data) described in Section 7 of the main paper, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets.
@@
The initial values for the VB algorithms are set as follows: the elements of the variational mean vector ($\boldsymbol{\mu}$) corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ are initialised using estimates obtained by fitting the SEM-Gau via maximum likelihood (ML) estimation. The variational means corresponding to the parameters $\nu$ and $\gamma$ are initialised to 4 and 1, respectively. For the models with Student’s $t$ errors, the set of latent variables $\boldsymbol{\tau}=(\tau_{1},\dots,\tau_{n})^{\top}$ must also be initialised. The variational means corresponding to these latent variables are initialised by sampling from the prior distribution $p(\tau_{i}\mid\nu^{(0)})\sim\text{IG}(\frac{\nu^{(0)}}{2},\frac{\nu^{(0)}}{2})$} for $i=1,\dots,n$, where $\nu^{(0)}=4$ denotes the initial value of $\nu$. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors.
[…]
S8.1 Results for the dataset 1
@@
Table S4 shows the posterior means and the 95% credible intervals for some of the model parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t obtained using the VB algorithm applied to the simulated dataset 1. The table also includes the computation times (in seconds) for one VB iteration, as well as the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values for each model, calculated using Equations (21) and (22), respectively, in Section 5 of the main paper.
Table S4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 (full data), along with the computation time (CT) in seconds for one VB iteration. The table also presents the DIC values for each model. Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.